OpenStudy (anonymous):

A container has 6 liters of a 40% alcohol solution in it. How much pure alcohol should be added to raise it to a 60% solution?

OpenStudy (amistre64):

6liters(40%) + yliters(100%) = (6+y)liters(60%) 6(.4) + y(1) = (6+y)(.6) 2.4 + y = 3.6 + .6y y-.6y = 3.6 - 2.4 y(1-.6) = 1.2 y(.4) = 1.2 y=(1.2)/(.4) = 12/4 = 3 6(.4) + 3(1) = 9(.6) 2.4 + 3 = 5.4 5.4 = 5.4 Yes!!!! Add 3 liters of 100% solution; but dont get it on your clothes or youll ruin them :)